OpenStudy (anonymous):

(3.6 × 109) ÷ (4 × 105)

OpenStudy (mathstudent55):

This is what you mean, right? \( \dfrac{3.6 \times 10^9}{4 \times 10^5}\)

OpenStudy (mathstudent55):

If so, then what is 3.6/4 = ? Also, what is \( \dfrac{10^9}{10^5} =\) ?